OpenStudy (anonymous):

solve by elimination or substitution 2x-4y=18 and 3x+5y=-3

hero (hero):

2x - 18 = 4y ---> y = (2x-18)/4 3x + 5y = -3 ---> y = (-3 - 3x)/5 y = y 5(2x - 18) = 4(-3 - 3x) 10x - 90 = -12 - 12x 10x + 12x = -12 + 90 22x = 78 x = 39/11

hero (hero):

y = -30/11

hero (hero):

you find y by plugging the x value back into one of the original equations

OpenStudy (anonymous):

thank you

hero (hero):

yw :D

OpenStudy (anonymous):

how did you go from 22x = 78 to x = 39/11

hero (hero):

x = 78/22 x = 39/11 reduction by 2

OpenStudy (anonymous):

oh, wow. great